# Bidiagonal matrix factorisations related to multiple orthogonal polynomials

### v1

@@ 4.1. Bidiagonal factorisations and connection with branched continued fractions
(Euler-Gauss method for branched continued fractions, cf. [20, Prop. 2.3]) For $r\in\mathbb{Z}^{+}$, let $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ be a sequence of formal power series with constant term $1$ and $f_{k}(t)={g_{k+1}(t)}/{g_{k}(t)}$ for all $k\in\mathbb{N}$. Then, the following conditions are equivalent:
  1. $\bullet$
    $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies the recurrence relation (4.3),
  2. $\bullet$
@@
    $\left(f_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies the relations in (4.7), so $f_{k}(t)$ is the generating function for $r$-Dyck paths at level $k$ with weights $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ and admits the Stieltjes-type $r$-branched-continued-fraction representation (4.8),
  3. $\bullet$
@@
Accordingly to [20, Prop. 8.2], the Hessenberg matrix $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$, involving the bidiagonal matrices in (1.2), is the production matrix of the matrix of generalised $r$-Stieltjes-Rogers polynomials $\left(S_{n,k}^{\,(r)}\left(\bm{\alpha}\right)\right)_{n,k\in\mathbb{N}}$. Therefore, based on [17, Thm. 4.2 & Prop. 4.4], the polynomial sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ satisfying the recurrence relation (4.1) is $r$-orthogonal with respect to $\left(v_{1},\cdots,v_{r}\right)$ defined by
$v_{j}\left[x^{n}\right]=S_{n}^{\,(r;\,j-1)}\left(\bm{\alpha}\right)\quad\text{for all }1\leq j\leq r.$ (4.10)
In particular, $v_{j}\left[1\right]=S_{0}^{\,(r;\,j-1)}\left(\bm{\alpha}\right)=1$ for all $1\leq j\leq r$.
Let $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ be the sequence of formal power series defined by (4.9), with $g_{0}(t)=1$. From the definition, it is clear that $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies (4.2). Using Lemma 4.2, $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ also satisfies (4.3) and is the unique solution of (4.3) such that ${g_{0}(t)=1}$. Finally, (4.2) determines uniquely the system $\left(v_{1},\cdots,v_{r}\right)$. ∎
@@
Let $\left(v_{1},\cdots,v_{r}\right)$ be a system of linear functionals such that there exists a sequence of formal power series $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ satisfying (4.2) and (4.3). Then, the nontrivial entries of the bidiagonal matrices appearing in the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of the Hessenberg recurrence matrix of the system $\left(v_{1},\cdots,v_{r}\right)$ are the coefficients of the branched continued fraction (4.8) for $f_{k}(t)={g_{k+1}(t)}/{g_{k}(t)}$ for all $k\in\mathbb{N}$.
Moreover, the formulas in (3.43) were obtained in [1, Cor. 4], within a completely combinatorial context where the $\left(\alpha_{k}\right)_{k\in\mathbb{N}}$ are the coefficients of a branched continued fraction equivalent to the one in (4.8). Therefore, although the bidiagonal factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of a Hessenberg matrix is not unique when $r\geq 2$, the factorisations obtained for a system $\left(v_{1},\cdots,v_{r}\right)$ via the methods developed in this section and in the previous section are the same.
4.2. A fundamental counterexample: symmetric multiple orthogonal polynomials
As mentioned in the introduction, bidiagonal matrix factorisations are not possible for all systems of multiple orthogonal polynomials. A counterexample is obtained when we consider $(r+1)$-fold symmetric $r$-orthogonal polynomials. We highlight it here, because it plays an important role in Subsection 4.3.
For $r\geq 1$, a polynomial sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ is $(r+1)$-fold symmetric if
$P_{n}\left(\mathrm{e}^{\frac{2\pi i}{r+1}}x\right)=\mathrm{e}^{\frac{2n\pi i}{r+1}}P_{n}(x)\quad\text{for all }n\in\mathbb{N}.$ (4.11)
The definition (4.11) is equivalent to the existence of $r+1$ polynomial sequences $\left(P_{n}^{[j]}(x)\right)_{n\in\mathbb{N}}$, $0\leq j\leq r$, such that
$P_{(r+1)n+j}(x)=x^{j}\,P^{[j]}_{n}\left(x^{r+1}\right)\quad\text{for all }n\in\mathbb{N}\text{ and }j\in\mathbb{N}.$ (4.12)
We refer to the sequences $\left(P_{n}^{[j]}(x)\right)_{n\in\mathbb{N}}$ as the components of the $(r+1)$-fold decomposition of $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$.
@@ 4.2. A fundamental counterexample: symmetric multiple orthogonal polynomials
$P_{n+r+1}(x)=x\,P_{n+1}(x)-\gamma_{n}\,P_{n}(x)\quad\text{for all }n\in\mathbb{N},$ (4.13)
with $P_{j}(x)=x^{j}$ for $0\leq j\leq r$ and $\gamma_{n}\neq 0$ for all $n\in\mathbb{N}$. When $r=1$, $(r+1)$-fold symmetric $r$-orthogonal polynomials reduce to symmetric orthogonal polynomials. Well-known examples include the classical Hermite, Tchebyshev, Legendre, and Gegenbauer polynomials.
The Hessenberg matrix $\mathrm{H}$ associated with the $(r+1)$-fold symmetric $r$-orthogonal polynomial sequence satisfying the recurrence relation (4.13) only has nonzero entries in the supradiagonal, all equal to $1$, and in the $r^{\text{th}}$-subdiagonal, equal to the recurrence coefficients $\left(\gamma_{n}\right)_{n\in\mathbb{N}}$, i.e.,
@@
The spectral properties of this type of recurrence matrix were investigated in [3]. If the Hessenberg matrix $\mathrm{H}$ in (4.14) admits a factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ as a product of bidiagonal matrices of the form in (1.2), then $\alpha_{0}=0$ and $\alpha_{0}\alpha_{r}\cdots\alpha_{r^{2}}=\gamma_{0}\neq 0$, which is clearly impossible.
4.3. Conditions for existence of the factorisations
The following theorem is the main result of this section. Firstly, it gives necessary and sufficient conditions on a system of linear functionals to have a sequence of formal power series $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ satisfying (4.2) and (4.3) and, equivalently, for the Hessenberg recurrence matrix of the system to admit a factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$, where all nontrivial entries of the bidiagonal matrices are different from 00. Furthermore, it connects the bidiagonal factorisations related to branched continued fractions to the Darboux transformations appearing in Subsection 3.3 and to the $(r+1)$-fold symmetric multiple orthogonal polynomials introduced in Subsection 4.2 and it establishes a so far missing link between branched continued fractions and type I multiple orthogonal polynomials.
Theorem 4.4.
For $r\in\mathbb{Z}^{+}$, let $\left(v_{1},\cdots,v_{r}\right)$ be a system of linear functionals such that $v_{j}\left[1\right]=1$ for any $1\leq j\leq r$. Then, there exists a sequence of formal power series $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ with constant term $1$ satisfying (4.2) and (4.3), with $\alpha_{k}\neq 0$ for all $k\in\mathbb{N}$, if and only if there is existence and uniqueness of all the multiple orthogonal polynomials on the step-line with respect to the $r+1$ systems of linear functionals
$V^{[j]}=\left(v_{j+1},\cdots,v_{r},x\,v_{1},\cdots,x\,v_{j}\right)\quad\text{for}\quad 0\leq j\leq r.$ (4.15)
Moreover, when these conditions hold:
  1. $\bullet$
    the $(r+1)$-fold symmetric $r$-orthogonal polynomial sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ satisfying the recurrence relation
@@ 1. $\bullet$
    with initial conditions $P_{j}(x)=x^{j}$ for $0\leq j\leq r$ and the same nonzero coefficients $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ as in (4.3), admits the decomposition
    $P_{(r+1)n+j}(x)=x^{j}\,P^{[j]}_{n}\left(x^{r+1}\right)\quad\text{for all }n\in\mathbb{N}\text{ and }0\leq j\leq r,$ (4.17)
    where $\left(P_{n}^{[j]}(x)\right)_{n\in\mathbb{N}}$ is the $r$-orthogonal polynomial sequence with respect to the system $V^{[j]}$,
  2. $\bullet$
    the sequence $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ is given by
    $g_{0}(t)=1\quad\text{and}\quad g_{(r+1)m+j+1}(t)=\sum_{n=0}^{\infty}q_{m}^{[j]}\left[x^{n+m}\right]\,t^{n}\quad\text{for all }m\in\mathbb{N}\text{ and }0\leq j\leq r,$ (4.18)
    where $\left(q_{m}^{[j]}\right)_{m\in\mathbb{N}}$, $0\leq j\leq r$, is the sequence of type I linear forms on the step-line with respect to $V^{[j]}$.
Recalling (4.9), the formula (4.18) gives a combinatorial interpretation for the nontrivial moments of the type I linear forms on the step-line with respect to the systems $V^{[j]}$.
Corollary 4.5.
For all $0\leq j\leq r$, let $\left(q_{m}^{[j]}\right)_{m\in\mathbb{N}}$ be the sequence of type I linear forms on the step-line with respect to $V^{[j]}$. Then, $q_{m}^{[j]}\left[x^{n+m}\right]=S_{n}^{\,(r;\,(r+1)m+j)}\left(\bm{\alpha}\right)$ for all $n,m\in\mathbb{N}$ and $0\leq j\leq r$.
Proof of Theorem 4.4.
Firstly, we suppose that there exists a sequence of formal power series $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ satisfying (4.2) and (4.3). We let $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ be the polynomial sequence satisfying the recurrence relation (4.16), with initial conditions $P_{j}(x)=x^{j}$ for $0\leq j\leq r$ and the same nonzero coefficients $\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ as in (4.3), and $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$, $0\leq j\leq r$, be the components of the $(r+1)$-fold decomposition (4.17) of $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$.
Based on [16, Thm. 4.7], the sequences $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$ are $r$-orthogonal with respect to the systems $V^{[j]}$ for all $0\leq j\leq r$, which implies that we have existence and uniqueness of the multiple orthogonal polynomials on the step-line with respect to any system $V^{[j]}$.
@@ 2. $\bullet$
Reciprocally, we suppose that the multiple orthogonal polynomials on the step-line with respect to the systems $V^{[j]}$ exist and are unique for all $0\leq j\leq r$ and we let $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$ and $\left(q_{k}^{[j]}\right)_{k\in\mathbb{N}}$ be the sequences of type II multiple orthogonal polynomials and type I linear forms on the step-line with respect to $V^{[j]}$.
Let $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ be defined by (4.17), i.e, be the $(r+1)$-fold symmetric polynomial sequence whose components are the sequences $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$. Then, the $r$-orthogonality of all $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$ with respect to $V^{[j]}$ implies that $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ is $r$-orthogonal with respect to $\left(u_{1},\cdots,u_{r}\right)$ such that
$u_{i}\left[x^{(r+1)n+k}\right]=\begin{cases}v_{i}\left[x^{n}\right]&\text{if }k=i-1,\\ \hfil 0&\text{if }k\neq i-1,\end{cases}\quad\text{for all }1\leq j\leq r,\;0\leq k\leq r,\text{ and }n\in\mathbb{N}.$ (4.19)
The multiple orthogonality conditions of $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ with respect to $\left(u_{1},\cdots,u_{r}\right)$ can be checked directly using the multiple orthogonality of the sequences $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$ with respect to the systems $V^{[j]}$. The calculations are completely analogous to the ones in the proof of [16, Thm. 4.7], so we do not repeat them here.
Therefore, $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ is a $(r+1)$-fold symmetric $r$-orthogonal polynomial sequence. Hence, $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ satisfies a recurrence relation of the form (4.16) for some nonzero coefficients $\left(\alpha_{k}\right)_{k\in\mathbb{N}}$. We will show that $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ defined by (4.18) satisfies (4.2) and (4.3), with the same coefficients $\left(\alpha_{k}\right)_{k\in\mathbb{N}}$ as in (4.16).
By definition, $q_{0}^{[j]}=v_{j+1}$ for all $10\leq j\leq r-$ 1. Therefore, $\left(g_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies (4.2).
@@
Let $\left(q_{k}\right)_{k\in\mathbb{N}}$ be the sequence of type I linear forms on the step-line with respect to $\left(u_{1},\cdots,u_{r}\right)$, dual to the sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$. Accordingly to [16, Prop. 4.3], $q_{k}\left[x^{\ell}\right]=0$ unless $\ell=k+(r+1)n$ for some $n\in\mathbb{N}$, and
$q_{(r+1)m+j}\left[x^{(r+1)(n+m)+j}\right]=q_{m}^{[j]}\left[x^{n+m}\right]\quad\text{for all }n,m\in\mathbb{N}\text{ and }0\leq j\leq r.$ (4.20)
As a result, (4.18) is equivalent to
@@
Moreover, because $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ satisfies (4.16), $\left(q_{k}\right)_{k\in\mathbb{N}}$ satisfies the dual recurrence relation

### v2

@@ 4. Bidiagonal factorisations linked to branched continued fractions
(cf. [22, Prop. 2.3]) For $r\in\mathbb{Z}^{+}$, let $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ be a sequence of formal power series with constant term 11 and $f_{k}(t)={g_{k+1}(t)}/\penalty 50{g_{k}(t)}$ for all $k\in\mathbb{N}$. Then, the following conditions are equivalent:
  • $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfies the recurrence relation ((c)),
@@
  • $\left(f_{k}(t)\right)_{k\in\mathbb{N}}$ satisfies the relations in (2.22), so $f_{k}(t)$ is the generating function for $r$-Dyck paths at level $k$ with weights $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ and admits the Stieltjes-type $r$-branched-continued-fraction representation (2.23),
@@
Accordingly to [22, Prop. 8.2], the Hessenberg matrix $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ is the production matrix of the matrix of generalised $r$-Stieltjes-Rogers polynomials of type 0, $\left(S_{n,k}^{\,(r;\,0)}\left(\bm{\alpha}\right)\right)_{n,k\in\mathbb{N}}$. As a result, the equivalence between (a) and (b) is a direct consequence of [17, Thm. 4.2 & Prop. 4.4].
The equivalence between (b) and (c) follows from Lemma 4.2. The sequence $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ and the system $\left(v_{1},\cdots,v_{r}\right)$ in (c) are uniquely determined by (4.5), with $g_{0}(t)=1$, and (4.2), respectively. ∎
As a consequence of Theorem 4.1, we can construct bidiagonal matrix factorisations for Hessenberg matrices associated with multiple orthogonal polynomials by finding sequences of formal power series $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfying ((c))-((c)) for systems of linear functionals $\left(v_{1},\cdots,v_{r}\right)$.
Taking the derivative at $t=0$ on both sides of ((c)), we find that if a sequence of power series $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ with constant term $1$ satisfies ((c)), then
$\alpha_{m}=g_{m+1}^{\,\prime}(0)-g_{m}^{\,\prime}(0)\quad\text{for all }m\in\mathbb{N}.$ (4.6)
If $\alpha_{m}\neq 0$, we can invert ((c)) to obtain
$g_{m+r+1}(t)=\frac{1}{\alpha_{m}\,t}\Big(g_{m+1}(t)-g_{m}(t)\Big).$ (4.7)
Therefore, starting from a system of linear functionals $\left(v_{1},\cdots,v_{r}\right)$, with $v_{j}\left[1\right]=1$ for all $1\leq j\leq r$, we can construct the sequences $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ and $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ satisfying ((c))-((c)) by setting $g_{0}(t)=1$, defining $g_{1}(t),\cdots,g_{r}(t)$ by ((c)), and then recursively obtaining $\alpha_{m}$ and $g_{m+r+1}(t)$ for all $m\in\mathbb{N}$, using formulas (4.6) and (4.7), respectively, unless there exists $m\in\mathbb{N}$ such that $\alpha_{m}=0$. Then, using Theorem 4.1, the nonzero coefficients $\alpha_{k}$ obtained correspond to nontrivial entries of the bidiagonal matrices appearing in the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of the Hessenberg recurrence matrix of the type II multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$.
@@
Let $\left(v_{1},\cdots,v_{r}\right)$ be a system of linear functionals such that there exists a sequence of formal power series $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfying ((c))-((c)). Then, the nontrivial entries of the bidiagonal matrices appearing in the factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of the Hessenberg recurrence matrix for the system $\left(v_{1},\cdots,v_{r}\right)$ are the coefficients of the branched continued fraction (2.23) for $f_{k}(t)={g_{k+1}(t)}/\penalty 50{g_{k}(t)}$ for all $k\in\mathbb{N}$.
Moreover, the formulas in (3.21) were obtained in [1, Cor. 4], within a completely combinatorial context where the $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ are the coefficients of a branched continued fraction equivalent to the one in (2.23). Therefore, although the bidiagonal factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ of a Hessenberg matrix is not unique when $r\geq 2$, the factorisations obtained for a system $\left(v_{1},\cdots,v_{r}\right)$ via the methods developed in this section and in Section 3 are the same.
In the following result, we give a necessary and sufficient condition for the recurrence matrix of a polynomial sequence to admit a bidiagonal factorisation of the form (1.1).
Theorem 4.4.
A polynomial sequence $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ satisfies a recurrence relation of the form
$x\left[P_{n}^{[0]}(x)\right]_{n\in\mathbb{N}}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}\left[P_{n}^{[0]}(x)\right]_{n\in\mathbb{N}},$ (4.8)
involving bidiagonal matrices $\mathrm{L}_{1},\cdots,\mathrm{L}_{r}$, and $\mathrm{U}$ of the form in (1.2), with all nontrivial entries $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ different from zero, if and only if $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ is $r$-orthogonal with respect to a system of linear functionals $V^{[0]}=\left(v_{1},\cdots,v_{r}\right)$ such that the multiple orthogonal polynomials on the step-line with respect to all systems
$V^{[j]}=\left(v_{j+1},\cdots,v_{r},x\,v_{1},\cdots,x\,v_{j}\right),\quad 0\leq j\leq r,$ (4.9)
introduced in (3.1), exist and are unique.
To prove that the existence and uniqueness of the multiple orthogonal polynomials on the step-line with respect to all systems $V^{[j]}$, $0\leq j\leq r$, is a sufficient condition for $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ to satisfy a recurrence relation of the form (4.8), we need to recall the definition and basic properties of $(r+1)$-fold symmetric $r$-orthogonal polynomials.
For $r\geq 1$, a polynomial sequence $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ is $(r+1)$-fold symmetric if $P_{n}\left(\mathrm{e}^{\frac{2\pi i}{r+1}}x\right)=\mathrm{e}^{\frac{2n\pi i}{r+1}}P_{n}(x)$ for all $n\in\mathbb{N}$. This definition is equivalent to the existence of $r+1$ polynomial sequences $\left(P_{n}^{[j]}(x)\right)_{n\in\mathbb{N}}$, $0\leq j\leq r$, which we refer to the as the components of the $(r+1)$-fold decomposition of $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$, such that
$P_{(r+1)n+j}(x)=x^{j}\,P^{[j]}_{n}\left(x^{r+1}\right)\quad\text{for all }n\in\mathbb{N}\text{ and }0\leq j\leq r.$ (4.10)
@@
$P_{n+r+1}(x)=x\,P_{n+r}(x)-\gamma_{n}\,P_{n}(x)\quad\text{for all }n\in\mathbb{N},$ (4.11)
with initial conditions $P_{n}(x)=x^{n}$ for $0\leq n\leq r$ and $\gamma_{n}\neq 0$ for all $n\in\mathbb{N}$. When $r=1$, $(r+1)$-fold symmetric $r$-orthogonal polynomials reduce to symmetric orthogonal polynomials. Well-known examples include the classical Hermite, Tchebyshev, Legendre, and Gegenbauer polynomials.
Remark 4.5.
The Hessenberg recurrence matrix $\mathrm{H}$ of the $(r+1)$-fold symmetric $r$-orthogonal polynomial sequence satisfying the recurrence relation (4.11) only has nonzero entries in the supradiagonal, all equal to $1$, and in the $r^{\text{th}}$-subdiagonal, equal to the recurrence coefficients $\left(\gamma_{n}\right)_{n\in\mathbb{N}}$, i.e.,
@@
If the Hessenberg matrix in (4.12) admits a factorisation $\mathrm{H}=\mathrm{L}_{1}\cdots\mathrm{L}_{r}\,\mathrm{U}$ as a product of bidiagonal matrices of the form in (1.2), then $\alpha_{0}=0$ and $\alpha_{0}\alpha_{r}\cdots\alpha_{r^{2}}=\gamma_{0}\neq 0$, which is clearly impossible. Therefore, $(r+1)$-fold symmetric $r$-orthogonal polynomial sequences provide an example of Hessenberg recurrence matrices which do not admit a bidiagonal matrix factorisation as in (1.1). The spectral properties of Hessenberg matrices of the form in (4.12), with positive and bounded coefficients $\left(\gamma_{n}\right)_{n\in\mathbb{N}}$, were investigated in [4].
Proof of Theorem 4.4.
Suppose that $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ is $r$-orthogonal with respect to $V^{[0]}=\left(v_{1},\cdots,v_{r}\right)$ such that the multiple orthogonal polynomials on the step-line with respect to all systems $V^{[j]}$, $0\leq j\leq r$, in (4.9) exist and are unique. Recalling Theorem 3.1, we know that the Hessenberg recurrence matrices associated with the systems $V^{[j]}$ admit the bidiagonal factorisations in (3.2), with $\alpha_{m}\neq 0$ for all $m\in\mathbb{N}$. In particular, $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ satisfies (4.8).
Reciprocally, suppose that $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ satisfies (4.8), with $\alpha_{m}\neq 0$ for all $m\in\mathbb{N}$. Recalling Theorem 4.1, $\left(P_{n}^{[0]}(x)\right)_{n\in\mathbb{N}}$ is $r$-orthogonal with respect to a system of linear functionals $\left(v_{1},\cdots,v_{r}\right)$ for which there exists a sequence of formal power series $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfying ((c))-((c)) with the same nonzero coefficients $\left(\alpha_{m}\right)_{m\in\mathbb{N}}$ appearing in (4.8). Let $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ be the $(r+1)$-fold symmetric $r$-orthogonal polynomial sequence satisfying the recurrence relation
@@
with initial conditions $P_{j}(x)=x^{j}$ for $0\leq j\leq r$, and let $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$, $0\leq j\leq r$, be the components of the $(r+1)$-fold decomposition (4.10) of $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$. Accordingly to [18, Thm. 4.7], the sequences $\left(P_{n}^{[j]}\right)_{n\in\mathbb{N}}$ are $r$-orthogonal with respect to the systems $V^{[j]}$ for all $0\leq j\leq r$. As a result, we know that the multiple orthogonal polynomials on the step-line with respect to all these systems exist and are unique. ∎
@@
In the following result, we explicitly describe the sequence $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ introduced in Theorem 4.1, using the moments of the type I linear forms on the step-line with respect to the systems $V^{[j]}$ in (4.9).
Proposition 4.6.
For $r\in\mathbb{Z}^{+}$ and a system $\left(v_{1},\cdots,v_{r}\right)$ of linear functionals, with $v_{j}\left[1\right]=1$ for all $1\leq j\leq r$, for which the multiple orthogonal polynomials on the step-line with respect to all systems $V^{[j]}$ in (4.9) exist and are unique, let $\left(q_{k}^{[j]}\right)_{k\in\mathbb{N}}$, $0\leq j\leq r$, be the sequences of type I linear forms on the step-line with respect to the systems $V^{[j]}$. Then, the sequence of formal power series $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfying ((c))-((c)) is given by
$g_{0}(t)=1\quad\text{and}\quad g_{(r+1)k+j+1}(t)=\sum_{n=0}^{\infty}q_{k}^{[j]}\left[x^{k+n}\right]\,t^{n}\quad\text{for all }k\in\mathbb{N}\text{ and }0\leq j\leq r.$ (4.14)
As mentioned in the introduction, when there exist a measure $\mu$ and weight functions $w_{1}(x),\cdots,w_{r}(x)$ defined in the support of $\mu$ such that $v_{j}\left[f\right]=\int f(x)w_{j}(x)\mathrm{d}\mu(x)$ for any $f\in\mathbb{C}[x]$ and any $1\leq j\leq r$, we can define the type I functions $Q_{k}^{[j]}(x)=\sum_{j=1}^{r}A_{\vec{n},j}(x)\,w_{j}(x)$ and the type I linear forms $q_{k}^{[j]}$ have the integral representation $q_{k}^{[j]}\left[f\right]=\int f(x)Q_{k}^{[j]}(x)\mathrm{d}\mu(x)$ for any $f\in\mathbb{C}[x]$. In that case, (4.14) implies the following integral representation for the functions $g_{m}(t)$, $m\geq 1$, using the Stieltjes transform of the type I functions $Q_{k}^{[j]}(x)$, valid for all $k\in\mathbb{N}$ and $0\leq j\leq r$:
$\frac{1}{z}\,g_{(r+1)k+j+1}\left(\frac{1}{z}\right)=\sum_{n=0}^{\infty}\frac{1}{z^{n+1}}\,\int x^{k+n}\,Q_{k}^{[j]}(x)\mathrm{d}\mu(x)=\int\frac{x^{k}\,Q_{k}^{[j]}(x)\mathrm{d}\mu_{j}(x)}{z-x}.$ (4.15)
We can use Proposition 4.6 to obtain a combinatorial interpretation for the moments of the type I linear forms on the step-line with respect to the systems $V^{[j]}$ as generating polynomials of partial $r$-Dyck paths, establishing a link of type I multiple orthogonal polynomials with lattice paths and branched continued fractions. Comparing (4.14) and (4.5), we find that
$q_{k}^{[j]}\left[x^{k+n}\right]=S_{n}^{\,(r;\,(r+1)k+j)}\left(\bm{\alpha}\right)=S_{n+k,k}^{\,(r;\,j)}\left(\bm{\alpha}\right)\quad\text{for all }k,n\in\mathbb{N}\text{ and }0\leq j\leq r.$ (4.16)
The second equality holds because, by definition of the modified and generalised $r$-Stieltjes-Rogers polynomials, both $S_{n}^{\,(r;\,(r+1)k+j)}\left(\bm{\alpha}\right)$ and $S_{n+k,k}^{\,(r;\,j)}\left(\bm{\alpha}\right)$ are both equal to the generating polynomial of the partial $r$-Dyck paths from $(0,0)$ to $\big((r+1)(n+k)+j,(r+1)k+j\big)$. In addition, we have $q_{k}^{[j]}\left[x^{m}\right]=0=S_{m,k}^{\,(r;\,j)}\left(\bm{\alpha}\right)$ whenever $m<k$. Therefore, we deduce that
$q_{k}^{[j]}\left[x^{m}\right]=S_{m,k}^{\,(r;\,j)}\left(\bm{\alpha}\right)\quad\text{for all }k,m\in\mathbb{N}\text{ and }0\leq j\leq r.$ (4.17)
This formula was given in [18, Prop. 4.3] without any mention of type I multiple orthogonality (in [18], the sequences $\left(q_{k}^{[j]}\right)_{k\in\mathbb{N}}$, $0\leq j\leq r$, appear as the dual sequences of the $r$-orthogonal polynomials with respect to the systems $V^{[j]}$).
Proof of Proposition 4.6.
It is trivial that $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ defined by (4.14) satisfies ((c)).
By definition, $q_{0}^{[j]}=v_{j+1}$ for all $0\leq j\leq r-1$. As a result, $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ also satisfies ((c)).
Now we need to prove that $\left(g_{m}(t)\right)_{m\in\mathbb{N}}$ satisfies ((c)). Let $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ be the $(r+1)$-fold symmetric $r$-orthogonal polynomial sequence satisfying the recurrence relation (4.13) and $\left(q_{k}\right)_{k\in\mathbb{N}}$ be the sequence of type I linear forms on the step-line dual to $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$. Accordingly to [18, Prop. 4.3]:
  • $q_{k}\left[x^{n}\right]=0$ if nkn-k is not a multiple of $r+1$,
  • $q_{(r+1)m+j}\left[x^{(r+1)k+j}\right]=q_{m}^{[j]}\left[x^{k}\right]=S_{m,k}^{\,(r;\,j)}\left(\bm{\alpha}\right)$ for all $m,k\in\mathbb{N}$ and $0\leq j\leq r$.
Therefore, (4.14) is equivalent to
@@
The recurrence relation (4.13) satisfied by $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ can be expressed as $x\left[P_{n}(x)\right]_{n\in\mathbb{N}}=\mathrm{H}\left[P_{n}(x)\right]_{n\in\mathbb{N}}$, where $\mathrm{H}$ is the Hessenberg matrix in (4.12) with $\gamma_{n}=\alpha_{n}$ for all $n\in\mathbb{N}$. Then, $\left(q_{k}\right)_{k\in\mathbb{N}}$ satisfies the dual recurrence relation $x\left[q_{k}\right]_{k\in\mathbb{N}}=\mathrm{H}^{\top}\left[q_{k}\right]_{k\in\mathbb{N}}$, which can be explicitly written as